OpenStudy (anonymous):

solve using the elimination method 9x + 11y = 86 9x – 2y = -5 please if you understand this method explain it because I do not understand how to use it

OpenStudy (asnaseer):

ok - you have two equations and have two unknowns to find - x and y the process involves using one equation to eliminate one of the variables. so, from the first equation we have: 9x + 11y = 86 therefore 9x = 86 - 11y now substitute this into the second equation to get: 9x - 2y = -5 therefore 86 - 11y - 2y = -5 therefore 86 - 13y = -5

OpenStudy (asnaseer):

from the last equation you should be able to work out what y is. then use this value of y in any of the original equations to work out what x is.

OpenStudy (asnaseer):

does that make sense?

OpenStudy (anonymous):

so you eliminated 9x all together?

OpenStudy (asnaseer):

yes

OpenStudy (anonymous):

ok that is what I was doing wrong thank you so much

OpenStudy (asnaseer):

yw

OpenStudy (anonymous):

So if im correct there is no solution?

OpenStudy (asnaseer):

there is a solution. the last equation we got to was: 86 - 13y = -5 therefore 13y = 86 + 5 = 91 therefore y = 7 now use the equation 9x + 11y = 86 to work out the value for x.